The points that passes through the line (2,3) and is perpendicular to the line that passes through the points (4,5) and (9,12) ..?
Answer by Cromlix(4381)

Hi there,
Gradient (slope) of line that passes through
(4,5) and (9,12)
Gradient (slope) = y2 - y1/x2 - x1
Gradient (slope) = 12 -5/9 - 4
Gradient (slope) = 7/5
Lines that are perpendicular to one another
have gradients (slopes) that multiply together
to give -1
m1 x m2 = -1
Using 7/5
7/5 x m2 = -1
m2 = -5/7
Setting up the line using the line equation
y - b = m(x - a)
m = -5/7 (a,b) = (2,3)
y - 3 = -5/7(x - 3)
y - 3 = -5/7x + 15/7
y = -5/7x + 15/7 + 21/7 (3)
y = -5/7x + 36/7
or multiply through by 7
7y = -5x + 36.
Points on this line can be:-
(3,3)
7y = -5(3) + 36
7y = -15 + 36
7y = 21
y = 3
(10,-2)
7y = -5x + 36
7y = -5(10) + 36
7y = -14
y = -2.
Hope this helps :-)